Shift report 20020106 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Jan 06 2002 - 16:03:28 EST

  • Next message: Apache: "Shift report 20020107 16:00-24:00"

    ------------------------------------------------------------
    Date: Sun Jan  6 16:03:28 2002, Shift: 08:00-16:00
    Supervisor: mads, with: 
    Log book#12 pages 8 - 9
    
    Shift Summary:
    ------------------------------------------------------------
    Magnet fields in D1-D4 were set to 0
    Polarity changed from A to B for magnets D1-D4
    
    During the shift there was no long period with 
    contious beam; The beam was ramped and dumped 
    many times.
    
    >From 1440 to 1500 there was a cogged beam. At 
    this time TPM1 tripped, and this had to be 
    fixed. Due to the short duration of the beam 
    and because TPM1 tripped, no data runs were taken.
    
    Took pedestal runs 6307-6310.
    
    Bravo still gives error message "can't connect
    to mainframe" when checking H1, H2 and C1
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Jan 06 2002 - 16:04:02 EST